# Observing network dynamics through sentinel nodes

### v1

@@ Abstract
A fundamental premise of statistical physics is that the particles in a physical system are interchangeable, and hence the state of each specific component is representative of the system as a whole. This assumption breaks down for complex networks, in which nodes may be extremely diverse, and no single component can truly represent the state of the entire system. It seems, therefore, that to observe the dynamics of social, biological or technological networks, one must extract the dynamic states of a large number of nodes—a task that is often practically prohibitive. To overcome this challenge, we use machine learning techniques to detect the network’s sentinel nodes, a set of network components whose combined states can help approximate the average dynamics of the entire network. The method allows us to assess the state of a large complex system by tracking just a small number of carefully selected nodes. The resulting sentinel node set offers a natural probe by which to practically observe complex network dynamics.
@@ I Introduction
We, therefore, seek efficient reduction methods, that help track the state of the system in a more compact fashion. In most applications, such reduction is achieved by focusing on the average activity of all nodes, i.e. $\langle x\rangle=\sum_{i=1}^{N}a_{i}x_{i}$, where the weights $\{a_{1},\ldots,a_{N}\}$ characterize the importance of each node in assessing the system’s dynamic state [9, 10, 11, 12, 13, 14, 15]. For example, selecting $a_{i}=k_{i}$, a node’s weighted degree, leads to the Gao-Barzel-Barabási (GBB) reduction [9], which helps predict the critical points of transition between dynamic states. Alternatively, under the dynamics approximate reduction technique (DART), the weight $a_{i}$ represents the $i$th entry of the leading eigenvector of the network’s adjacency matrix [10, 11].
While these reduction methods provide crucial theoretical insight, allowing us to analytically predict the system’s critical points of transition, they offer limited advances on observability. This is because measuring $\langle x\rangle$ via GBB or DART, one still needs to access the state of all nodes $x_{1},\dots,x_{N}$, thus remaining beyond the practical bounds of empirical accessibility. More selective probes have been introduced in the context of early warning signals, where one follows just a small number of nodes that show signs of looming transitions early on, prior to the rest of the network [16, 17, 18, 19, 20]. The problem is that such early warning nodes are, by design, selected thanks to their divergent behavior around the transition points. This renders them unlikely to be a representative sample of the entire networked system.
How then can we gain the predictive power of GBB or DART, but with a node set that is comparable to the few early warning signifiers? To solve this, here we use machine learning to impose a sparsity condition on $\{a_{1},\ldots,a_{N}\}$, requiring that $a_{i}\neq 0$ only for a small number of nodes. The result is a set of sentinel nodes, whose average allows us to approximate the state of the entire system. We find that, with only a few strategically selected nodes, we can achieve network-wide observability.
Our analysis shows that the ideal sentinel node sets tend to sample the network heterogeneity, for example, comprising a range of low, average and high degrees. This indicates that the optimal combination of sentinel nodes is mainly determined by the network topology, largely independent of the system’s specific dynamics. Consequently, one can extract the sentinels even without a priori knowledge about the system’s hidden dynamics. This allows practical observability even under extremely restrictive conditions.
@@ II.1 Seeking the sentinel nodes
Consider coupled nonlinear dynamics on a network, such as mutualistic or epidemic dynamics. (See Methods for the dynamics models and networks used.) In many cases, we want to know the mere average of the network activity, i.e., $\overline{x}=\sum_{i=1}^{N}x_{i}/N$, where $x_{i}$ is the activity of the $i$th node, and $N$ is the number of nodes in the network. For example, in contagion processes, if we assign $x_{i}=0$ for susceptible and $x_{i}=1$ for infectious, then the average $\overline{x}$ represents the fraction of infectious individuals in the population. Similarly, $\overline{x}$ may represent the average political opinion in a population, the fraction of people aware of news, average biomass in an ecosystem, or the extent of cancerous cells in the body. Quantity $\overline{x}$ is simple, intuitive, and practical, providing a first-hand summary of dynamics on networks.
Denoting by $x_{i}^{*}$ the equilibrium of $x_{i}$, our goal is to approximate the average activity of all nodes, i.e.,
$\overline{x}=\frac{1}{N}\sum_{i=1}^{N}x_{i}^{*},$ (1)
using a limited set $S$ of sentinel nodes, namely,
$\overline{x}^{\prime}=\frac{1}{n}\sum_{i\in S}x_{i}^{*}.$ (2)
Here $n\ll N$ is the number of sentinel nodes in $S$. To obtain $S$, we simulate the system under a range of conditions, e.g., by varying the average edge weight $D$. Parameter $D$ is a natural control parameter; as $D$ varies, $x$ in Eq. (1) may change or undergo critical transitions (Fig. 1). We then seek the optimal node set $S$ that minimizes
@@
where $\overline{x}_{\ell}$ and $\overline{x}^{\prime}_{\ell}$ are $\overline{x}$ and $\overline{x}^{\prime}$, respectively, at the $\ell$th value of $D$. Equation (3) captures the total difference between the approximation $\overline{x}^{\prime}(D)$ and the exact $\overline{x}(D)$ across the $L$ different conditions that together sample a range of $D$ values. It is minimal when $\overline{x}^{\prime}$ successfully recovers the system average in Eq. (1), including its observed points of transition.
As an example, we consider the coupled double-well dynamics (Fig. 1a) on the dolphin social network (Fig. 1b) under different coupling strength $D$. When uncoupled, each node has two states, i.e., lower and upper states, as equilibria. We initialize each simulation for each $D$ value by placing all the nodes in their lower state. We show the equilibrium state values, $x_{i}^{*}$, for all nodes as a function of $D$ by the light gray lines in Fig. 1c–h. The black line represents the average activity of the network, $\overline{x}$. As $D$ increases, $\overline{x}$ increases in stages with a prolonged intermediate range in which some $x_{i}^{*}$s remain in their lower states (i.e., $x_{i}^{*}<2$) and the other $x_{i}^{*}$ have transitioned to their upper states (i.e., $x_{i}^{*}>5$). This range of $D$ values captures a multistage transition [21, 22, 19].
Figure 1: Approximations of the average activity, $\overline{x}$, of the coupled double-well dynamics on the dolphin network. (a) Coupled double-well dynamics model. (b) Dolphin network with $N=62$ nodes. Larger circles marked with color represent the nodes used in sentinel node approximations shown in (c)–(f): orange for $n=1$, yellow for $n=2$, green for $n=3$, and blue for $n=4$. (c)–(h) Equilibrium states in the dolphin network. The gray lines indicate $x_{i}^{*}$ at each value of the control parameter. The black lines indicate $\overline{x}$. Note that the gray and black lines are the same in panels (c)–(h). The purple lines in (c)–(f) show the sentinel node approximation corresponding to the selected sentinel nodes, whose $x_{i}^{*}$ values are shown in orange (c), yellow (d), green (e), and blue (f). We verified by exhaustive search of every possible combination of nodes that the obtained node sets for $n=1$ (shown in (c)) and $n=2$ (shown in (d)) are exact minimizers of $\varepsilon$. Panels (g) and (h) show the GBB and DART approximations in red and pink, respectively. The brown and dark yellow lines in each plot show the observables (i.e., a particular weighted average of $\{x_{1}^{*},\ldots,x_{N}^{*}\}$) that the GBB and DART schemes intend to approximate.
We demonstrate approximating $\overline{x}$ with the sentinel node approximation, $\overline{x}^{\prime}$, with $n=1$, $2$, $3$, and $4$ by the purple line in Fig. 1c, 1d, 1e, and 1f, respectively. The sentinel node approximation correctly identifies the onset of the bifurcation at $D\approx 0.2$ for any $n\in\{1,\ldots,4\}$ but over- or underestimates $\overline{x}$ depending on $D$ in the range of the multistage transition. In general, larger optimized node sets do not include smaller ones as a subset. In this example, there is no overlap between any of the optimized node sets with $n=1,2,3$, and $4$; we show the selected sentinel nodes in color in Fig. 1b. However, the approximation accuracy progressively improves as one includes more sentinel nodes. Concretely, the approximation errors for $n=1$, $2$, $3$, and $4$ are $\varepsilon=0.035$, $\varepsilon=0.020$, $\varepsilon=0.016$, and $\varepsilon=0.004$, respectively. Figure 1f demonstrates that, with a microscopic number of nodes (i.e., $n=4$), the sentinel node approximation captures major components of the multistage transition, resulting in a small approximation error. As expected, further increasing $n$ results in further reduction in approximation error (see SI section S2); however, for all practical purposes $n=4$ is already sufficient.
For comparison, we show the results obtained from the GBB and DART approximations in Fig. 1g and 1h, respectively. Even with just $n=1$ sentinel node, our method performs better than the GBB and DART, reducing $\varepsilon$ by more than 75% (GBB: $\varepsilon=0.171$; DART: $\varepsilon=0.142$). This is mainly owing to the fact that, while GBB and DART track the collective state of the system (i.e., $\langle x\rangle$ in the Introduction section) and thus predict a single transition point, our method can discern the multistage nature of the actual observed transition (black solid line). Most crucially, thanks to its machine learning component, our framework achieves the observed predictability with a much smaller set of observations—here just with $n=4$ nodes (Fig. 1f). In sum, our method realizes a dramatic reduction in complexity along with a comparable, and in the present case even improved, accuracy.
Because our algorithm is stochastic, we ran it 100 times, starting with independent initial node sets, on the same network to assess generality of the results shown in Fig. 1. We set $n=4$. We show the approximation error, $\varepsilon$, for the 100 optimized node sets by the blue circles in Fig. 2a. Each circle corresponds to one run. The optimized node set is different in each run in general. However, the $\varepsilon$ values of these optimized node sets are similar to each other, and most importantly they are small. This result supports the robustness of our algorithm with respect to the initial condition and stochasticity of the algorithm.
Figure 2: Approximation error for the coupled double-well dynamics for node sets with $n=\lfloor\ln N\rfloor$ nodes. A blue circle represents an optimized node set. An orange triangle represents a degree-preserving random node set. A green square represents a completely random node set. (a) Dolphin network. (b) BA network with $N=10^{3}$ nodes.
Next, we select two groups of random node sets for comparison. First, we take the best optimized node set among the 100 runs in terms of $\varepsilon$, denoted by $\tilde{S}$. We then make 100 node sets by randomly selecting $n=4$ nodes having the same degree sequence as $\tilde{S}$. For example, if the selected optimized node set is composed of four nodes with degrees 1, 3, 3, and 8, then we sample one node with degree 1 uniformly at random from the network and similarly sample two nodes with degree 3 and one node with degree 8. We call this node set the degree-preserving random node set. We show $\varepsilon$ for each degree-preserving random node set by the orange triangles in Fig. 2a. Second, we make 100 node sets with $n=4$ by selecting nodes uniformly at random from the entire network. We call this node set the completely random node set. We calculate $\varepsilon$ for each of the 100 completely random node sets, shown by the green squares in Fig. 2a. As expected, the degree-preserving node sets outperform the completely random ones, exhibiting an approximation error that is concentrated on smaller values in general. The crucial point is, however, that our optimized node sets have consistently smaller errors than both alternatives, better than 100% of the random sets, and 89% of the degree-preserving sets. Therefore, while selecting $n$ nodes with appropriate degrees improves the accuracy of approximation, one can substantially reduce $\varepsilon$ by further optimization beyond only using a good degree sequence.
These results remain similar for other networks; see Fig. 2b for results for a Barabási-Albert (BA) network and SI section S3 for results obtained from eight more networks. In all networks, the optimized node sets yield relatively small $\varepsilon$.
The patterns described for the coupled double-well dynamics across the various networks also hold true for the mutualistic species, gene-regulatory, and SIS dynamics (see SI section S3). Across every combination of one of the four dynamics and one of the ten networks, the worst optimized node set still has a smaller $\varepsilon$ than 100% of completely random node sets and 95.2% of degree-preserving node sets.
@@
We statistically verified these observations by constructing an analysis of variance (ANOVA) model of the approximation error, $\varepsilon$. The dependent variable is $\ln\varepsilon$. There are three qualitative independent variables: dynamics (reference: coupled double-well dynamics), network (reference: BA network), and node set type (reference: completely random node sets). We exclude the Erdős-Rényi (ER) network because it yields small $\varepsilon$ regardless of node set type (see SI section S3), presumably because of the small spread of its degree distribution and the relative similarity across all nodes. We show the details of our analysis in SI section S4. Across all dynamics and networks, $\varepsilon$ of the optimized node set is on average more than 200 times smaller than $\varepsilon$ of completely random node sets ($p<10^{-7}$) and more than 25 times smaller than $\varepsilon$ of degree-preserving node sets ($p<10^{-7}$).
While we have used undirected and unweighted networks, we have verified that these results, including the ANOVA results, remain similar for various weighted networks (see SI section S9) and directed networks (see SI section S10).
II.2 Characteristics of good node sets
We have shown that finding nodes with an appropriate degree sequence is an important but not sufficient condition for finding a good set of sentinel nodes, $S$. In this section, we explore the characteristics of the nodes that our algorithm selects for inclusion in $S$.

### v3

@@ Abstract
A fundamental premise of statistical physics is that the particles in a physical system are interchangeable, and hence the state of each specific component is representative of the system as a whole. This assumption breaks down for complex networks, in which nodes may be extremely diverse, and no single component can truly represent the state of the entire system. It seems, therefore, that to observe the dynamics of social, biological or technological networks, one must extract the dynamic states of a large number of nodes—a task that is often practically prohibitive. Theoretical tools are also highly restrictive, given the analytically impenetrable combination of complex heterogeneous networks with nonlinear, often hidden, dynamics. To overcome this challenge, we use machine learning techniques to detect the network’s sentinel nodes, a set of network components whose combined states can help approximate the average dynamics of the entire network. The method allows us to assess the equilibrium state of a large complex system by tracking just a small number of carefully selected nodes. We find that the sentinels are mainly determined by the network structure such that they can be extracted even with little knowledge of the system’s specific interaction dynamics. Therefore, the network’s sentinels offer a natural probe by which to observe the system’s dynamic states. Intriguingly, sentinels tend to avoid the highly central nodes such as the hubs.
@@ I Introduction
We, therefore, seek efficient reduction methods, that help track the state of the system in a more compact fashion. In most applications, such reduction is achieved by focusing on the average activity of all nodes, i.e., $\langle x\rangle=\sum_{i=1}^{N}a_{i}x_{i}$, where the weights $\{a_{1},\ldots,a_{N}\}$ characterize the importance of each node in assessing the system’s dynamic state [9, 10, 11, 12, 13, 14, 15]. For example, selecting $a_{i}=k_{i}$, a node’s weighted degree, leads to the Gao-Barzel-Barabási (GBB) reduction [9], which helps predict the critical points of transition between dynamic states. Alternatively, under the dynamics approximate reduction technique (DART), the weight $a_{i}$ represents the $i$th entry of the leading eigenvector of the network’s adjacency matrix [10, 11].
While these reduction methods provide crucial theoretical insight, allowing us to analytically predict the system’s critical points of transition, they offer limited advances on observability. This is because measuring $\langle x\rangle$ via GBB or DART, one still needs to access the state of all nodes $x_{1},\dots,x_{N}$, often beyond the bounds of empirical accessibility. More selective probes have been introduced in the context of early warning signals, where one follows just a small number of nodes that show signs of looming transitions early on, prior to the rest of the network [16, 17, 18, 19, 20]. The problem is that such early warning nodes are, by design, selected thanks to their divergent behavior around the transition points. This renders them unlikely to be a representative sample of the entire networked system.
How then can we capture the system’s collective dynamics, but with a node set that is comparable to the few early warning signifiers? To solve this, here we use machine learning to impose a sparsity condition on $\{a_{1},\ldots,a_{N}\}$, requiring that $a_{i}\neq 0$ only for a small number of nodes. The result is a set of sentinel nodes, whose average allows us to approximate the equilibrium state of the entire system. We find that, with only a few strategically selected nodes, we can achieve network-wide observability. The challenge is conceptually related to the functional observability problem in control theory [21, 22]. The crucial distinction, however, is that our computational approach is not confined to linear dynamics. Instead, it enables observation of nonlinear systems and their critical transitions, even in cases where the underlying nonlinear mechanisms are unknown.
Our analysis shows that an ideal sentinel node set tends to sample the network heterogeneity. For instance, it typically comprises a range of low, average, and high degree nodes. Therefore, the optimal combination of sentinel nodes is mainly determined by the network topology, largely independent of the system’s specific dynamics. Consequently, one can extract the sentinels even without a priori knowledge about the system’s hidden dynamics. This result allows practical observability even under restriction that one lacks full knowledge of the system’s internal driving dynamic mechanisms.
@@ II.1 Seeking the sentinel nodes
Consider a network with $N$ components (nodes), whose activities $x_{i}(t)$ are driven by nonlinear dynamics. This can capture, e.g., mutualistic dynamics in ecological networks, epidemic dynamics on social systems, or the double-well dynamics presented in Fig. 1a. We denote the stable equilibrium states of the system by $x_{i}^{*}$. While the full equilibrium state of the system comprises all node activities $x_{i}^{*}$, $i=1,\dots,N$, quite often, we only seek the average network activity
$\overline{x}=\frac{1}{N}\sum_{i=1}^{N}x_{i}^{*}$ (1)
to help characterize the system’s global state. For example, in contagion processes, assigning $x_{i}=0$ for susceptible and $x_{i}=1$ for infectious, the average $\overline{x}$ represents the fraction of infectious individuals in the population at equilibrium. In a similar fashion, $\overline{x}$ may represent the mean political opinion, the fraction of people aware of news, the average biomass in an ecosystem, or the proliferation of cancerous cells in the body, all depending on the scientific context. Quantity $\overline{x}$ is simple, intuitive, and practical, providing a natural parametrization of the network’s dynamics.
Our goal is to approximate $\overline{x}$ in Eq. (1) using a limited set $S$ of sentinel nodes, namely,
$\overline{x}^{\prime}=\frac{1}{n}\sum_{i\in S}x_{i}^{*},$ (2)
where $n\ll N$ is the number of sentinel nodes in $S$. To obtain $S$, we simulate the system under a range of conditions, e.g., by varying the average edge weight $D$ in Fig. 1a. As $D$ is varied, $\overline{x}$ changes accordingly, and at specific points, it may also undergo a critical transition. A good sentinel node set $S$, then, satisfies two conditions: (i) It optimally approximates the observed $\overline{x}$ as $D$ varies, including its critical transitions; (ii) It achieves this with a small number of sentinels $n$.
Therefore, we simulate the system under $L$ different conditions $D=D_{1},\dots,D_{L}$, and seek the optimal node set $S$ that minimizes
@@
where $\overline{x}_{\ell}$ and $\overline{x}^{\prime}_{\ell}$ are $\overline{x}$ and $\overline{x}^{\prime}$ as obtained under $D_{\ell}$, respectively. Equation (3) captures the total difference between the approximation $\overline{x}^{\prime}(D)$ and the exact $\overline{x}(D)$ across the $L$ different conditions that sample the examined range of $D$ values. It is minimal when $\overline{x}^{\prime}$ successfully recovers the complete system average of Eq. (1), including its observed points of transition.
As an example, we consider the coupled double-well dynamics (Fig. 1a) implemented on a dolphin social network, displayed in Fig. 1b. We first examine the full system average $\overline{x}$ vs. $D$ (Fig. 1c-f, black solid lines). Here, the uncoupled system ($D=0$) has two equilibrium points—one high and the other low. In our simulations we initiated the system’s state around the low equilibrium, and let it relax from there. For small $D$, the system settles at that low state. However, as $D$ is increased, $\overline{x}$ undergoes a critical transition, favoring the high equilibrium point. This transition begins at around $D=0.2$, but it is not completely abrupt. Instead, it features discrete steps, starting at $D\approx 0.2$, and only reaching a full transition at $D\approx 0.6$. This is because the individual node states $x_{i}^{*}$ (grey solid lines) do not all transition at the exact same point; some shift early on, other only at a later stage. Therefore, we observe a multistage transition [23, 24, 19].
Figure 1: Observing network dynamics via sentinel nodes. Approximations of the average activity, $\overline{x}$, of the coupled double-well dynamics on the dolphin network. (a) Coupled double-well dynamics. (b) Dolphin network with $N=62$ nodes. Larger circles (colored) represent nodes in four different sentinel node sets $S$ with sizes $n=1$ (orange), $n=2$ (yellow), $n=3$ (green), and $n=4$ (blue). (c)–(f) The equilibrium states $x_{i}^{*}$ of all nodes in the dolphin network as a function of $D$, obtained from numerical simulations (grey lines). The average state $\overline{x}$ is also shown (black solid lines). The purple solid lines represent the sentinel node approximations $\overline{x}^{\prime}$, obtained from the four node sets highlighted in (b). For $n=1$, the approximation substantially deviates from the actual $\overline{x}$ (yielding $\varepsilon=0.035$; see panel (c)) because, just a single sentinel node is used. This deviation is gradually reduced in (d), (e), and (f) as $n$ is increased, until it is practically eliminated at $n=4$ ($\varepsilon=0.004$). In (c)–(f), we also show the individual sentinel node states (red, yellow, green and blue lines).
Next, we seek to approximate the $\overline{x}$ vs. $D$ curve by tracking a small number $n$ of sentinel nodes determined by our optimization algorithm. In Fig. 1c, we show the outcome for $n=1$ (orange line). To be specific, we selected the best single sentinel out of the $N$ nodes, which minimizes $\varepsilon$ in Eq. (3). We observe a gap between the sentinel approximation (purple) and the exact $\overline{x}$ (black), and hence we continue to search for larger sentinel node sets $S$. Indeed, as we increase $n$ in Fig. 1d-f, we find that the sentinel approximation follows the black solid line more closely. At $n=4$, we find that $S$ successfully recovers the observed multistage transition, yielding $\varepsilon=0.004$, a negligible discrepancy from the full system average. Notably, the smaller sets are not included as part of the larger ones, namely, the single node ($n=1$) in Fig. 1c is not part of the sentinel node sets with $n=2$ (Fig. 1d), $3$ (Fig. 1e), or $4$ (Fig. 1f). The sentinel property is a characteristic of the set $S$, not of its individual elements.
Figure 1 demonstrates our key result. It shows that we can approximate the average of the full state of the system by tracking just a microscopic subset of $n=4$ nodes. This tiny sentinel node set, $S$, captures the equilibrium state of the system under the entire range of $D$ values, including the critical transition and its multistage transition pattern. This result represents a crucial advance, since it is often extremely difficult to probe the entire system, and hence being able to track $\overline{x}$ via such a small $S$ is highly desirable. Increasing $n$ results in a reduced $\varepsilon$, thus improving the approximation even further. However, for all practical purposes, we find that there is no need to go beyond $n=4$.
For illustration purposes, our demonstration has focused on a relatively small network ($N=62$) coupled with a specific form of nonlinear dynamics (i.e., double-well). To complement this, in Supplementary Section S1, we detail our expanded testing ground, comprising a broad set of $20$ model and empirical networks, whose sizes range from $N=62$ to $N=81,171$ nodes. We test each of these networks with four nonlinear dynamics, i.e., coupled double-well, ecological, epidemic, and gene-regulatory. In Supplementary Section S2 (see Fig. S1), we examine the complete array of bifurcation curves $\overline{x}$ vs. $D$, covering all networks and dynamics. Similar to the results obtained in Fig. 1, we find that our optimized sets consistently recover the full system average activity, its bifurcations, and multistage transition patterns. In Supplementary Section S3, we also include a detailed comparison of our method against the theoretically motivated GBB and DART, alongside an additional GBB variant [12].
The process for detecting the optimal set $S$ is non-deterministic. Therefore, depending on the initial seed of $n$ nodes, the algorithm may settle on distinct final sets. To examine the consistency of the method’s predictive power, we ran our optimization $100$ times, starting from a different selection of nodes each time. As expected, each iteration led to a different final set of nodes. The crucial point is that, despite these differences, all the detected sentinel node sets exhibited good performance. For example, on the dolphin network, our sentinel node sets all yielded an error narrowly distributed around $\varepsilon\approx 5\times 10^{-3}$ (Fig. 2a, top panel, blue circles)
Figure 2: Sentinel node approximation for different networks and dynamics. (a) Approximation errors, $\varepsilon$, for 100 optimized (blue circles), degree-preserving (orange triangles), and completely random (green squares) node sets using the coupled double-well dynamics on 15 empirical and 5 model networks. The networks range in size from $N=62$ (Dolphin network) to $N=81,171$ (Prosper network); the model networks each have approximately $1,000$ nodes. ER: Erdős-Rényi. BA: Barabási-Albert. HK: Holme-Kim. GKK: Goh-Kahng-Kim. LFR: Lancichinetti-Fortunato-Radicchi. (b) Approximation errors for 100 node sets of each type using the SIS dynamics on the same 20 networks.
We compare the $100$ sentinel node sets with two natural alternatives. First, we examine completely random node sets, obtained by selecting $n$ nodes with uniform probability from across the network. Such random $n$-plets show highly inconsistent results, with $\varepsilon$ ranging across two or more orders of magnitude (Fig. 2a, green squares). These large and diverse errors were consistently observed across our $20$ networks. Hence, not surprisingly, our optimized $S$, perform substantially better than random. As a more challenging contender, we next consider degree-preserving random node sets (Fig. 2a, orange triangles). To construct these sets, we begin with the best performing optimized node set among the $100$ runs, namely, the one with the smallest $\varepsilon$. We denote this winning node set by $\tilde{S}$. A degree-preserving set consists of randomly selected nodes that have the same degree sequence as $\tilde{S}$. The idea is to generate node sets that have similar characteristics to the best performing set detected across all optimization runs. For example, if the $n=4$ nodes in $\tilde{S}$ have degrees $k=1,3,3$, and $8$, then we generate degree-preserving sets by sampling one node with degree $k=1$ uniformly at random from the network, two nodes with degree $k=3$, and one final node with degree $k=8$. Such degree-preserving node sets are superior to random sets, but are still consistently outperformed by our optimized $S$. Taken together, in Fig. 2a, we find that the optimized sets (blue) are better than $99.3\%$ of the degree-preserving node sets, and $100\%$ of the completely random node sets.
In Fig. 2b, we show a similar analysis for the susceptible-infectious-susceptible (SIS) epidemic dynamics, confirming similar results. This is further supported in Supplementary Section S4 (see Fig. S5), where we repeat this examination on two other dynamics models. Quite consistently, across all network/dynamics combinations, the optimized $S$ supersedes completely random and degree-preserving random node sets $100\%$ and $96.8\%$ of the time, respectively.
@@
In Supplementary Section S4 (see Table S1), we further verify our observations statistically through an analysis of variance (ANOVA) model of the approximation error $\varepsilon$. In this analysis, we set the dependent variable to be $\ln\varepsilon$, and considered three qualitative independent variables: dynamics (reference: coupled double-well dynamics), network (reference: BA network), and node set type (reference: completely random node sets). We exclude the Erdős-Rényi (ER) network because it yields small $\varepsilon$ in a majority of cases, regardless of the node set. This is presumably because of the relative similarity across all nodes characteristic in this family of networks (see Figs. 2 and S5). Across all pairs of network and dynamics, we observe that $\varepsilon$ of the optimized node set is, on average, more than 900 times smaller than $\varepsilon$ of the completely random sets ($p<10^{-7}$) and more than 40 times smaller than $\varepsilon$ of the degree-preserving sets ($p<10^{-7}$). Finally, in Supplementary Sections S5 and S6, we validate our algorithm’s performance also on weighted and directed networks. Overall, we find that our optimized sentinel node sets can uncover the global network behavior with just a handful of nodes, over a highly diverse corpus of network and dynamics, covering different scales and distinct scientific domains.
II.2 Characteristics of good individual sentinel nodes
Our analyses in Fig. 2 and in Supplementary Section S4 (see Table S1) have shown that finding nodes with an appropriate degree sequence is an important but not sufficient condition for finding a good set $S$ of sentinel nodes. In this section, we explore further the characteristics of the nodes that our algorithm selects for inclusion in $S$. To achieve this, in Supplementary Section S7, we use a random forest model to examine an array of node characteristics that can potentially correlate with the sentinel property (see Table S10). The first, most natural, characteristic we focus on is the node degree $k_{i}$. We ask whether sentinel nodes tend to have a distinguishable degree centrality. We find, however, that a node’s degree is a poor predictor of sentinelity. Nodes selected by our algorithm may be of low or intermediate degree, showing no specific tendency towards the average $\overline{k}$, as one may expect, given that they are selected to predict $\overline{x}$. Interestingly, we find that the largest hubs are significantly under-represented among the sentinel nodes (Supplementary Section S7; see Fig. S14). Consistent with this observation, sentinel nodes do not particularly tend to be outliers in terms of the $x_{i}^{*}$ values, either (see Supplementary Section S8). The sentinel node sets tend to avoid the hubs, focusing primarily on small and intermediate degree nodes.
Beyond degrees, we analyze the nodes’ nearest neighbor degree, the nodes’ local clustering, closeness centrality, betweenness centrality, and coreness. Our findings indicate that these quantities have relatively weak predictive power and cannot be reliably used to a priori identify potential sentinel nodes. Hence, there is little that can be said about the properties of nodes in $S$; they seem to have potentially diverse characteristics.
The reason for this, we believe, is that sentinelity is not an intrinsic property of individual nodes, but rather a characteristic of the set $S$. In other words, sentinelity emerges from the collective nature of all nodes in $S$, rather than from any specific component. Therefore, below, instead of focusing on individual sentinel nodes, we shift our attention to characterizing sentinel node sets.
Figure 3: Characteristics of sentinel nodes. (a) Degree histogram of the dolphin network. (b) Degree histogram extracted from $100$ single sentinel nodes ($n=1$). (c)–(e) Similar histograms across $100$ sets with $n=2$, 33, and 44 nodes. We find that sentinel node sets tend to spread across the degree sequence. Here, the bar color represents the node rank in $S$ in terms of degree: blue - smallest-degree node, orange - second smallest, red - third, light blue - fourth. (f) We repeated the analysis for a BA network of $N=10^{3}$ nodes. For this network, we observe the hub nodes at the tail of the histogram, with degrees ranging from $k=20$ to $85$ (inset). (g)–(j) Histograms extracted from the sentinel node sets with $n=1,2,3$, and $6$. All histograms with $n>2$ comprise a mixture of small and intermediate degree nodes, but lack any representation of the hub nodes (see empty insets). Bar colors are the same as in panels (b)–(e), with two additional ranks: green - fifth smallest node, yellow - sixth smallest node.
II.3 Characteristics of good sentinel node sets
We have shown that single-node features do not predict whether nodes are good sentinels or not. This is because a sentinel node set is a combinatorial construct, and not an independent collection of single nodes. Therefore, we now focus on the set features that may help characterize good sentinel node sets.